Calculate the volume in liters of a
×9.810−7/mmolL
mercury(I) chloride solution that contains
400.μmol
of mercury(I) chloride
Hg2Cl2
1
Expert's answer
2017-10-29T11:50:06-0400
400 μmol = 400x10-6 mol = 400x10-3 mmol = 0.4 mmol
c=n/V
V=n/c
V=(0.4 mmol)/(9.8×〖10〗^(-9) mmol/L)=4.1×〖10〗^7 L